OpenStudy (anonymous):

Use the quadratic formula to solve 2y2 – 3y = 2. A. {½, 2} B. {½, –2} C. {–½, –2} D. {–½, 2}

pooja195 (pooja195):

Before the subtract the 2 so the the equation is = to 0 \[\huge 2y^2-3y-2\]

pooja195 (pooja195):

What are your a b c values?

pooja195 (pooja195):

Once you have those we can plug it into this equation AKA Quadratic Formula

pooja195 (pooja195):

\[\huge~x=\frac{ -b \pm \sqrt{b^2-4ac} }{ 2a }\]

OpenStudy (anonymous):

I barley started to learn this, do you think you can walk me through it, step by step?

pooja195 (pooja195):

\[\huge~ax^2+bx+c\] \[\huge~2y^2-3y-2=0\] Anything that is ^2 will be the A just tell the number ignore the y^2 then anything with a plain vairble will just be the b value in this case it is -3 A number alone is the C value which in this case is -2

pooja195 (pooja195):

Got it so far?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

@pooja195

pooja195 (pooja195):

Good now we plug it in

OpenStudy (kendricklamar2014):

Step 1. Subtract 2 from both sides: \[2y^2−3y−2=2−2\] \[= 2y^2−3y−2=0\]

OpenStudy (anonymous):

okay! @KendrickLamar2014

OpenStudy (kendricklamar2014):

Step 2. Factor the Left side: \[(2y+1)(y−2)=0\]

pooja195 (pooja195):

\[\huge~x=\frac{ -(-3)\pm \sqrt{(-3)^2-4(2)(-2)} }{ 2(2) } \]

OpenStudy (kendricklamar2014):

Step 3. Set the factors to equal 0: \[2y+1=0~ or ~y−2=0\]

pooja195 (pooja195):

@KendrickLamar2014 it says to use the quadratic formula not factoring

OpenStudy (kendricklamar2014):

I know, but you can get the same answer.

pooja195 (pooja195):

They need to use the quadratic formula its the directions.

OpenStudy (kendricklamar2014):

\[= y = -\frac{ 1 }{ 2 }~or~ y = 2\]

pooja195 (pooja195):

Or just give them the answer...... @ninowletonzalez06 are you allowed to use factoring?

OpenStudy (kendricklamar2014):

Yah, I was just showing her that there are many ways to solve this equation

pooja195 (pooja195):

Or have u not learned it because the Quadratic formula comes before factoring @KendrickLamar2014

pooja195 (pooja195):

Well now i bet that the asker is confused .-.

OpenStudy (anonymous):

a liitle, can you explain it in the easiest form!

pooja195 (pooja195):

do you HAVE to use the quadratic formula?

OpenStudy (anonymous):

by what the question is asking yes!

OpenStudy (anonymous):

are you still there?

pooja195 (pooja195):

Ok :) All i did here was plug in the values

pooja195 (pooja195):

LaTex isnt working i cant finish it .-.

pooja195 (pooja195):

\[\huge~x=\frac{ -(-3) \pm \sqrt{(-3)^2-4(2)(-2)} }{ 2(2)} \]

OpenStudy (anonymous):

okay!:)

pooja195 (pooja195):

\[\huge~x=\frac{ 3 \pm \sqrt{(-3)^2-4(2)(-2)} }{ 2(2)} \] simplify whats in the square root but DO NOT square root it

OpenStudy (anonymous):